1. A 6.0 kg mass is connected over a pulley to a 3.0 kg mass. What is the resulting acceleration of the masses when they are released? What is the tension in the string when the masses are released?


1
Expert's answer
2021-07-20T09:46:18-0400

"m_1 = 6kg"

"m_2=3kg"

"\text{Projection of forces onto the vertical axis:}"

"-m_1a=-m_1g+T"

"m_2a=-m_2g+T"

"\text{Solution of this system of equations:}"

"a = g\frac{m_1-m_2}{m_1+m_2}=9.8*\frac{6-3}{6+3}=3.27\frac{m}{s^2}"

"T=\frac{2m_1m_2g}{m_1+m_2}=\frac{6*3*9.8}{6+3}=19.6N"

"\text{Answer: } a = 3.27\frac{m}{s^2};T=19.6N"